Part A Calculate the molar concentration of OH ions in a 0.555 M solution of hypobromite ion (BrO; Kb 4.0 x 10-) Express theWhat is the pH of this solution? Express the pH to two decimal places. VAD ΑΣφ ? pH=
Answer #1

A)

BrO- dissociates as:

BrO- +H2O -----> HBrO + OH-

0.555 0 0

0.555-x x x

Kb = [HBrO][OH-]/[BrO-]

Kb = x*x/(c-x)

Assuming x can be ignored as compared to c

So, above expression becomes

Kb = x*x/(c)

so, x = sqrt (Kb*c)

x = sqrt ((4*10^-6)*0.555) = 1.49*10^-3

since c is much greater than x, our assumption is correct

so, x = 1.49*10^-3 M

So, [OH-] = x = 1.49*10^-3 M

Answer: 1.5*10^-3 M

B)

use:

pOH = -log [OH-]

= -log (1.49*10^-3)

= 2.8268

use:

PH = 14 - pOH

= 14 - 2.8268

= 11.1732

Answer: 11.17